# Beyond Standard MoE: Mixture of Latent Experts for Resource-Efficient Language Models

### v1

@@ Appendix C Critical Role of the “Down Operator”
This section presents empirical evidence demonstrating that the “down operator” encapsulates more essential information than other components in Mixture of Experts (MoE) Large Language Models (LLMs). Through systematic experimentation, we establish the importance of preserving this operator’s structure for maintaining model performance.
@@
Consistent with our methodology in Appendix B, we utilize the Qwen1.5-MoE-A2.7B model with a fixed latent parameter $k=10$. To isolate the significance of the “down operator,” we implement two distinct transformation approaches:
@@
    Partial transformation (“up+gate”) - Converts only the “up operator” and “gate operator” to their Mixture of Latent Experts (MoLE) equivalents while preserving the original “down operator.”
@@
We evaluate these transformations across multiple downstream tasks, with results summarized in Table 4.
Transformation Type ($k=10$) GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext-2 PPL $\downarrow$
Original MoE 60.1 61.0 9.49
Partial (“up+gate”) 59.4 60.0 9.74
Complete (“all”) 43.2 57.0 10.63
Table 4: Performance comparison of Qwen1.5-MoE-A2.7B under different transformation configurations. Results demonstrate the critical importance of preserving the “down operator” structure for maintaining model performance.
The results in Table 4 reveal a clear performance hierarchy. The original MoE model achieves optimal performance across all metrics. The partial transformation, which preserves the “down operator,” exhibits only marginal performance degradation (0.7% decrease on GSM8K, 1.0% on MMLU, and 0.25 increase in perplexity). In stark contrast, the complete transformation that modifies all operators, including the “down operator,” suffers substantial performance deterioration (16.9% decrease on GSM8K, 4.0% on MMLU, and 1.14 increase in perplexity).
These findings provide compelling evidence that the “down operator” contains critical information that significantly impacts model performance. When this operator is transformed into the latent space, substantial information loss occurs, resulting in markedly diminished capabilities across reasoning and knowledge-intensive tasks. This asymmetric importance among the operators suggests that architectural modifications to MoE models should prioritize preserving the structure of the “down operator” to maintain performance integrity.
Appendix D Downstream Task Evaluation
To rigorously assess the generalization capabilities of the proposed MoLE architecture, we conducted comprehensive evaluations across a diverse set of downstream tasks. The comparative performance metrics are presented in Table 5.
Model Wikitext-2
(PPL $\downarrow$)
Standard MoE 75.86
MoLE 81.57
Table 5: Performance comparison between standard MoE and MoLE models across language modeling benchmark tasks. Bold values indicate superior performance.
The empirical results in Table 5 demonstrate that despite utilizing approximately 40% fewer parameters than the standard MoE architecture, the proposed MoLE model achieves comparable performance across the PPL tasks. This parameter efficiency without substantial performance degradation highlights the effectiveness of our proposed architectural modification.
Appendix E Training Arguments
Hyperparameters MoE MoLE
@@ Appendix E Training Arguments
Table 6: Model architecture and training hyperparameter configurations for MoE and MoLE models.

### v2

@@ Appendix B Critical Role of the “Up Operator”
While empirical experiments on Qwen1.5-MoE-A2.7B demonstrate that transferring all operators for MoE models can achieve superior performance, we observe that different operators contribute differentially to MoLAE transformation efficacy. This section presents empirical evidence establishing that the “up operator” encapsulates more essential information than other components in MoE models. Through systematic experimentation, we demonstrate the importance of preserving this operator’s structure for maintaining model performance.
@@
We examine a different MoE architecture, Moonlight-16B-A3B [Liu et al., 2025], in which the critical role of the “up operator” is more pronounced. Consistent with our methodology in Section 6.1, we utilize the Moonlight-16B-A3B model with a fixed latent parameter $k=8$, as it contains $64$ experts per layer. To isolate the significance of the “up operator,” we implement the following distinct transformation approaches:
@@
    Partial transformation (“up+gate”) - Converts only the “up operator” and “gate operator” to theirMoLAE equivalents while preserving the original “down operator.”
@@
    Partial transformation (“up+down”) - Converts only the “up operator” and “down operator” to theirMoLAE equivalents while preserving the original “gate operator.”
  3. 3.
    Partial transformation (“gate+down”) - Converts only the “gate operator” and “down operator” to theirMoLAE equivalents while preserving the original “up operator.”
  4. 4.
@@
We evaluate these transformations on the MMLU and CEval [Huang et al., 2023] tasks, with results summarized in Table 4.
Table 4: Performance comparison of Moonlight-16B-A3B under different transformation configurations. Results demonstrate the critical importance of preserving the “up operator” structure for maintaining model performance.
Task ($k=8$) Original “up+gate” “up+down” “gate+down” “all”
MMLU (%) $\uparrow$ 67.2 55.7 53.1 60.2 42.1
CEval (%) $\uparrow$ 74.2 57.6 52.9 62.8 46.6
In contrast to Qwen-MoE, the Moonlight-MoE architecture exhibits more substantial performance degradation under MoLAE transformation. As evidenced in Table 4, preserving the “up operator” yields optimal performance on downstream tasks. This phenomenon demonstrates the disproportionate importance of the “up operator” relative to the other components.
These findings provide compelling evidence that the “up operator” encodes critical information that significantly influences model performance. When this operator is transformed into the latent space, substantial information loss occurs, resulting in markedly diminished capabilities across reasoning and knowledge-intensive tasks. This asymmetric importance among operators suggests that architectural modifications to MoE models should prioritize preserving the structure of the “up operator” to maintain performance integrity.
Appendix C Training Arguments
See Table 5.
Table 5: Model architecture and training hyperparameter configurations for MoE and MoLAE models.
Hyperparameters MoE MoLAE
@@ Appendix C Training Arguments
Appendix D Refined MoLAE Transformation
In Section 5, we introduced the methodology to transform a standard MoE model into its corresponding MoLAE formulation. This section presents a refined approximation approach that incorporates activation information, resulting in enhanced precision.
Given activation matrices $X^{i}$ for $i\in\{1,2,\ldots,N\}$, our objective is to determine low-rank factorization matrices $A^{i}$ and $B$ such that $W^{i}X^{i}\approx A^{i}BX^{i}$. While Section 5 assumed that the effect of activation matrices could be eliminated—simplifying the problem to Equation (6)—we now develop a more robust approximation that explicitly incorporates activation information.
D.1 Problem Formulation
We formulate the refined approximation as an optimization problem to find matrices $A^{i}$ and $B$ that minimize the sum of Frobenius norm differences between the original expert computations and their low-rank approximations:
$\min_{A^{i},B}\quad F(A^{1},\ldots,A^{N},B):=\frac{1}{2}\sum_{i=1}^{N}\|W^{i}X^{i}-A^{i}BX^{i}\|_{F}^{2}$ (14)
To facilitate the solution, we introduce block matrices $W$, $A$, and $X$ defined as:
$W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix},\quad A=\begin{pmatrix}A^{1}\\ A^{2}\\ \vdots\\ A^{N}\end{pmatrix},\quad X=\text{diag}(X^{1},X^{2},\ldots,X^{N})$ (15)
This block representation transforms the problem in Equation (14) into an equivalent matrix factorization problem:
$\min_{A,B}\quad\frac{1}{2}\|WX-ABX\|_{F}^{2}$ (16)
This formulation enables us to derive an activation-aware low-rank approximation that more accurately preserves the input-output relationships of the original expert modules compared to the activation-agnostic approach described in Section 5.
Problem (16) cannot be solved directly using the Eckart-Young-Mirsky theorem due to the presence of the activation matrix $X$. We therefore establish the following theorem that characterizes an optimal solution to problem (16).
Theorem 2
Let $X$ be the activation matrix and $W$ be the weight matrix. Assume that $X^{\top}X$ is positive definite with Cholesky decomposition $X^{\top}X=LL^{\top}$ where $L$ is invertible. Let $L^{\top}W=U\Sigma V^{\top}$ be the singular value decomposition of $L^{\top}W$ and $\Sigma_{m}$ be the truncated diagonal matrix containing the $m$ largest singular values, with corresponding truncated matrices $U_{m}$ and $V_{m}$. Then $A^{*}=(L^{\top})^{-1}U_{m}\Sigma_{m}^{1/2}$ and $B^{*}=\Sigma_{m}^{1/2}V_{m}^{\top}$ constitute an optimal solution to problem (16).
Proof 2
Since $X^{\top}X$ is positive definite, there exists a unique Cholesky decomposition $X^{\top}X=LL^{\top}$ where $L$ is invertible. Utilizing this decomposition, we can transform the original optimization problem as follows:
$\|WX-ABX\|_{F}^{2}$ $=\text{Tr}[(WX-ABX)^{\top}(WX-ABX)]$
$=\text{Tr}[(X^{\top}(W-AB)^{\top}(W-AB)X]$
$=\text{Tr}[(W-AB)^{\top}XX^{\top}(W-AB)]$
$=\text{Tr}[(W-AB)^{\top}LL^{\top}(W-AB)]$
$=\text{Tr}[(L^{\top}(W-AB))^{\top}(L^{\top}(W-AB))]$
$=\|L^{\top}W-L^{\top}AB\|_{F}^{2}$
$=\|\tilde{W}-L^{\top}AB\|_{F}^{2},$
where $\tilde{W}=L^{\top}W$. Given that $\tilde{W}=U\Sigma V^{\top}$ is the SVD of $\tilde{W}$, by the Eckart-Young-Mirsky theorem, the best rank-$m$ approximation of $\tilde{W}$ is $\tilde{W}_{m}=U_{m}\Sigma_{m}V_{m}^{\top}$, which can be factorized as $\tilde{W}_{m}=\tilde{A}\tilde{B}$ where $\tilde{A}=U_{m}\Sigma_{m}^{1/2}$ and $\tilde{B}=\Sigma_{m}^{1/2}V_{m}^{\top}$.
We now demonstrate that $A^{*}=(L^{\top})^{-1}\tilde{A}$ and $B^{*}=\tilde{B}$ constitute an optimal solution to the original problem. First, for any matrices $A$ and $B$ with appropriate dimensions:
$\min_{A,B}\|L^{\top}AB-\tilde{W}\|_{F}\leq\|L^{\top}A^{*}B^{*}-\tilde{W}\|_{F}=\|\tilde{A}\tilde{B}-\tilde{W}\|_{F}$ (17)
Conversely, since $\tilde{A}\tilde{B}$ is the optimal rank-$m$ approximation of $\tilde{W}$:
$\|\tilde{A}\tilde{B}-\tilde{W}\|_{F}=\min_{\text{rank}(T)\leq m}\|T-\tilde{W}\|_{F}\leq\|L^{\top}AB-\tilde{W}\|_{F},\quad\forall A,B$ (18)
The final inequality holds because $\text{rank}(L^{\top}AB)\leq\text{rank}(AB)\leq m$ for any feasible solution $(A,B)$.
Combining these inequalities:
$\min_{A,B}\|L^{\top}AB-\tilde{W}\|_{F}=\|L^{\top}A^{*}B^{*}-\tilde{W}\|_{F}$ (19)
Therefore, $(A^{*},B^{*})$ is an optimal solution to problem (16), which completes the proof.
Based on Theorem 2, we propose the refined Algorithm 2 as follows.
Algorithm 2 Refined MoLAE Transformation
1:Weight matrices $W^{1},W^{2},\ldots,W^{N}$, activation matrices $X^{1},X^{2},\ldots,X^{N}$, and target rank $m$
2:Low-rank factorization matrices $A^{1},A^{2},\ldots,A^{N}$ and $B$
3:Construct block matrices $W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix}$ and $X=\text{diag}(X^{1},X^{2},\ldots,X^{N})$
4:Compute $X^{\top}X$
5:if $X^{\top}X$ is singular then
6:    Apply regularization: $X^{\top}X\leftarrow X^{\top}X+\lambda I$ for a small $\lambda>0$
7:end if
8:Compute the Cholesky decomposition: $X^{\top}X=LL^{\top}$
9:Compute $\tilde{W}=L^{\top}W$
10:Perform SVD on $\tilde{W}$: $\tilde{W}=U\Sigma V^{\top}$
11:Extract the $m$ largest singular values and corresponding singular vectors:
12:$\Sigma_{m}$, $U_{m}$, and $V_{m}$
13:Compute $\tilde{A}=U_{m}\Sigma_{m}^{1/2}$
14:Compute $B=\Sigma_{m}^{1/2}V_{m}^{\top}$
15:Compute $A=(L^{\top})^{-1}\tilde{A}$
16:Extract blocks of $A$ to obtain $A^{1},A^{2},\ldots,A^{N}$
17:return $A^{1},A^{2},\ldots,A^{N},B$
Appendix E Limitations
We acknowledge certain limitations inherent to the present investigation. Constraints on available computational resources precluded evaluations of models at exceptionally large scales, notably DeepSeek-V3-671B and DeepSeek-R1-671B. Consequently, our empirical analysis was performed utilizing the Moonlight-16B-A3B model. This model employs an architectural design identical to that of DeepSeek-R1-671B, as elaborated upon in Appendix B.